OpenStudy (anonymous):

I am working using the grouping formula. Can it be applied to factorise this equation?

OpenStudy (anonymous):

n^3 + 3 n^2 + 2 n – 720 = 0 (n^3 + 3 n^2) + (2 n – 720) = 0 n^2(n + 3) + 2(n – 360) = 0